OpenStudy (anonymous):

Find the x-intercepts of the graph of the equation: y = x^2 - 3x + 2 hmm so i got x=2, and x=1 which doesn't make sense

ganeshie8 (ganeshie8):

why ?

OpenStudy (anonymous):

well when i plugged the numbers in it wasn't equivalent

OpenStudy (freckles):

\[2^2-3(2)+2=0 ? \\ \text{ well } 2^2-3(2)+2=4-6+2=-2+2=0 \\ \text{ so } 2^2-3(2)+2=0 \text{ is a true equation }\]

OpenStudy (freckles):

and you can check the other one in a similar way

OpenStudy (freckles):

you should get \[1^2-3(1)+2 \text{ is also } 0\]

OpenStudy (anonymous):

oh wait so i got it right....

OpenStudy (freckles):

yes the x-intercepts are (2,0) and (1,0)

OpenStudy (freckles):

you solved 0=x^2-3x+2 correctly

OpenStudy (anonymous):

OMG!!!! proud moment :P

OpenStudy (freckles):

you did good! :)

OpenStudy (anonymous):

thanks! :)